OpenStudy (richyw):

I should know how to do this by now, but how do I find the maximum and minimum of \[6\sin\left(\frac{\pi}{3}x\right)+2\sin\left(\pi x\right)\] on \(0\leq x \leq 3\)

OpenStudy (richyw):

\[2\pi\cos\left(\frac{\pi}{3}x\right)+2\pi\cos\left(\pi x\right) =0\]

OpenStudy (science0229):

First, substitute\[u=\frac{ \pi }{ 3 }x\]

OpenStudy (science0229):

So, \[\pi x=3u\]

OpenStudy (science0229):

Then divide both sides by 2pi

OpenStudy (science0229):

What did you get?

OpenStudy (science0229):

hello?